Constants PartA The tield just outside a 3.30 cm-radius metal ball Is 3.75x102 N/C and points toward the ball. What charge resides on the ball? Express your answer to three significant tigures and include the appropriate units. Q-Value Units Submit Incorrect; Try Again, 2 attempts remaining < Return to Assignment Provide Feedback
Answer #1

For the metal ball, electric field just outside the ball is given by the expression -

E = k*Q / R^2

Where -

k = Coulomb's constant = 9.0 x 10^9 N*m^2/C^2

R = 3.30 cm = 0.033 m

Q = charge on the ball

E = 3.75 x 10^2 N/C (Given)

Put the values -

3.75 x 10^2 N/C = (9.0 x 10^9 x Q) / (0.033)^2

=> Q = (3.75 x 10^2 x (0.033)^2) / (9.0 x 10^9) = 0.04537 x 10^-9 C = 4.54 x 10^-11 C (Answer)